A balloon contains 245.4 L of helium at 4.00ºC and 1.00 atm. What is the temperature (in ºC) of the gas if the volume has increased by 324.1 L at the same pressure?
1
Expert's answer
2018-12-07T01:49:10-0500

V1/T1 = V2/T2

T2 = T1*V2/V1 = 4°C*324.1L/245.4L = 5.3 °C